answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How do you find the probability of an event followed by another event?

If the events are independent then you can multiply the individual probabilities. But if they are not, you have to use conditional probabilities.

If Drawing 3 cards that all queen from a deck of cards?

The probability of drawing three queens from a deck of cards is (4 in 52) times (3 in 51) times (2 in 50), or 24 in 132,600 or 1 in 5,525.

How do you convert percentages into averages?

Are you talking about grades? There are several ways to do this. Ask the teacher how it is done.

What does it mean for probability to be exhaustive?

A set of events is said to be exhaustive if, between them, they cover all possible outcomes.

What is the probability of rolling 5 on regular six sided six?

If you roll the die often enough, the probability is 1 - a certainty.

On a single roll, the probability is 1/6.

If you roll the die often enough, the probability is 1 - a certainty.

On a single roll, the probability is 1/6.

If you roll the die often enough, the probability is 1 - a certainty.

On a single roll, the probability is 1/6.

If you roll the die often enough, the probability is 1 - a certainty.

On a single roll, the probability is 1/6.

Probability of dealing ace king and queen of spades in that order when dealing only 3 cards?

The probability of dealing the Ace of Spades from a normal 52 card deck is 1 in 52. The probability of dealing the King of Spades next is 1 in 51. The probability of dealing the Queen of Spades next is 1 in 50.

The probability of drawing those three cards in that order is the product of those probabilities, which is 1 in 132,600. This is the same as the number of permutations of N (52) things taken P (3) at a time, which is stated as N! - P! (52 * 51 * 50)

If you did not care what order the cards were dealt in, but still wanted to know the probability of getting the Ace, King, and Queen of Spades, then you would be talking about the combinations of N (52) things taken P (3) at a time, which is stated as (N! - P!) / (N - P)! (52 * 51 * 50 / 3 / 2 / 1). The probability in this case is 1 in 22,100.

What is the probability of having a girl or boy child?

Statistically speaking, about 50% for each, but in reality it will depend on the father's genetics and timing of conception, pH level of the mother's body, and maybe even some more stuff.

When two dice are rolled what is the probability that both will be the same number?

6 sides on each dice. Each number on the sides is different, this represents the 1 out of all 6 amounts shown. So this is what you'll do for each dice...

1/6 + 1/6 = 2/6 = 1/3

What does sub standard material mean?

It means that the material does not meet at least one of the requirements in terms of its quality.

How do you increase your chances of having a girl?

girls are actually easier to conceive than boys. girl sperm are generally stronger. sperm counts lessen in men with heat, so tight underwear for men, position; woman on top,

What are the uses of probability in the working world?

It is used in insurance and health care to look at how risky it is to offer a policy to a given person or group. In Las Vegas, it is used in all games. When looking at reactions to medications, people look at the odds that a given person will have an adverse outcome. There are many more, but this gives you an idea.

If an identical twin marries an identical twin what is the likelihood of them having twins?

the statistics of twins marrying twins having twin children is only a slight fractional increase in percentage of chance. it is almost negligable however the slight increase has lead many people to believe the common misconseption of twim marraige resulting in the birth of more twins alot.

What is the chance that a given event will occur usually expressed between the number 0 and 1 (will not the event occor)?

The chance is equal to the proportion of people who do not understand the basics of probability. 1 does not represent the chance that "will not the event occor" or, that the event will not occur.

Trending Questions
What is the probability of getting a sum of seven in rolling two dice? What is the weight of a deck of cards? Is ace before 2? What is a forecast or prediction of the probable course and outcome of a disorder? What is the probability of rolling a number higher than 4? What is the probability that a seven-digit phone number contains the number 7? When two balanced dice are rolledthere are 36 possible outcomes. Find the probability that the sum is a multiple of 3 or greater than 4? What is the ideal severity rate? What does estimated experimental probability mean please someone answer? What is the probability of getting a red card from a deck of cards? A class has 8 students who are to be assigned seating by a lot What is the probability that the students will be arranged in order shortest to tallest assume that no two students are the same height? Why 1 is one and 2 is two? A jar contains 8 red marbles 9 blue marbles and 6 green marbles Two marbles are picked at random What is the probability that one is green and one is blue? Is probably spelled probably or probobly? Probability of two tetrahedral dice rolled the sum on the upward faces is four? How is the Pythagorean theorem used to develop the equation of a circle in standard form? A coin is fipped and a number cube is rolled so what is the sample space of the experiment? Does saltfish heads resemble human heads? What is the expected sum ofthe numbers that appear when three fair dice are rolled? What is the probability of rolling a number less that 3 on a number cube?